OpenStudy (anonymous):

write the value of (8 / 27) to the negative power of a third

OpenStudy (debbieg):

So, what you really want is to find an x such that: \[\Large x^{-3}=\dfrac{8}{27}\] So one approach is to take each side of that equation to the reciprocal power... eg: \(\Large (x^{-3})^{-1/3}=\left(\dfrac{8}{27}\right)^{-1/3}\) Which gives you \(\Large (x^{-3})^{-1/3}=x\) on the left hand side, so evaluate the right hand side and that it what you need to take to the power -3 to get \(\Large \dfrac{8}{27}\)